A company make widgets. One of every 70 widgets is faulty and has to be fixed. The company makes a $4 profit for each widget sold, but loses $72 for each faulty widget they have to fix. Find the expected value of each widget.
Answer #1

since One of every 70 widgets is faulty and has to be fixed so

let P(faulty widget )= P(F) = 1/70

P( not faulty widget ) = P(NF) = 1-(1/70)= 69/70

let x is the company makes a $4 profit for each widget and y is loses $72 for each faulty widget

expected value of each widget = x *P(NF) + y*P(F) = 4*[69/70] + (-72)*[1/70] = 204/70= 2.9143

the expected value of each widget is 2.9143